OpenStudy (anonymous):

when do you know to square ratios? for example: the widths of two similar rectangles are 16cm and 14cm. what is the ratio of the area?

OpenStudy (anonymous):

areas vary with the square. for example if you have a square of side 2 its area is 4, and if have a square of side 2*5= 10 its area is 100

OpenStudy (anonymous):

10 is 5 times 2 and the area is 25 times 4

OpenStudy (anonymous):

so supposes these were just squares. one would have area \[16^2=256\] and the other would have area \[14^2=196\]

OpenStudy (anonymous):

the ratio of 16 to 14 is 8 to 7, \[\frac{16}{14}=\frac{8}{7}\] the ratio of 256 to 196 is 64 to 49 \[\frac{256}{196}=\frac{64}{49}\]